# SG-Reg: Generalizable and Efficient Scene Graph Registration

### v1

@@ I INTRODUCTION
Early works in semantic-aided registration explicitly maintain the semantic objects. To match re-observed semantic objects, they design hand-crafted descriptors and similarity metrics based on 2D image feature[9], projected bounding box [12, 8], quadratic representation[13], or graph topology[14, 10, 15]. Using the calculated similarity metric, they identify matched nodes by either setting a threshold or designing a set of verification rules. Although these works introduced pioneering SLAM frameworks that incorporate semantic landmarks, they often face corner cases in real-world experiments due to the noise commonly present in the reconstructed semantic objects.
@@ II-C Scene Graph Construction
Fusion++[9] 2018 Explicit $\times$ $\times$ 2D Points Mask R-CNN[21] -
@@
Ours 2024 Implicit Point Cloud 3D Points RAM-Ground-SAM NetVLAD+LightGlue
@@ II-D Semantic Data Association
Data association between semantic representations is the most important step to guarantee a successful registration. In semantic SLAM works, The step relies on explicit representations. Bowman et al. [12] represent each object using a bounding box. If two object nodes belong to the same semantic category and their intersection-over-union (IoU) exceeds a predefined threshold, they are considered a match. Building on Kimera, Hydra [15] constructs hand-crafted semantic descriptors by aggregating semantic histograms from nearby objects and computes node similarity using these descriptors. More recently, SlideSLAM [53] determines object similarity based on identical semantic categories and similar bounding box shapes. To incorporate topological information into semantic data association, X-View [14] introduces a random walk descriptor (RWD) that records semantic labels along walking routes. This descriptor is then converted into a topological representation. When calculating the similarity metric, X-View counts the number of rows with identical RWDs to determine a similarity score. To enhance the spatial descriptiveness of RWD, Lin et al. [10] integrate distance information into the arrangement of RWDs. Similarly, Liu et al. [55] utilize spatial priors to construct RWDs. To reduce false matches between nodes, some approaches [56] explicitly construct edge descriptors to verify geometric consistency among matched nodes. Along the same lines, Julia et al. [57] employ node triplets to validate the correctness of matched nodes.
@@
Inspired by learning-based graph matching models [58], we believe training a neural network to learn semantic data association is a promising direction. The latest version of Hydra [50] proposes encoding object nodes using a Graph Neural Network (GNN), though it omits shape features. SGAligner [16] is the first work, to our knowledge, to focus on learning scene graph matching. It encodes multiple modalities from semantic nodes, including semantic labels, center positions, and relationship labels. Subsequently, SG-PGM enhances this approach by incorporating shape features into scene graph learning. It samples points from each semantic node and aggregates them using GATv2 [25]. SGAligner and SG-PGM adopted the geometric-related layers from GeoTransformer, including its superpoint matching, point matching, and local-to-global registration layer. Both of them are trained and evaluated on the 3RScan dataset [29], requiring ground-truth annotations to construct their scene graph.
@@ IV-B Triplet-boosted Graph Neural Network (GNN)
where the length of triplet edges $|\mathbf{e}_{i,j}|,|\mathbf{e}_{i,k}|\in\mathbb{R}$ and $\hat{\mathbf{e}}_{i,j},\hat{\mathbf{e}}_{i,k}$ are the normalized vectors. Sinusoidal functions $\psi^{L}$ and $\psi^{A}$ are applied to embed the edge length and the triplet angle. The sinusoidal function has been used in GeoTransformer[41] to encode their geometric scalars.
@@ IV-C Shape Network
In the following steps, we explain the encoding for $\{{}^{l}\mathcal{X}\}_{l}$ and it goes the same way for $\{{}^{l}\mathcal{Y}\}_{l}$. As shown in Fig. 4, the point backbone generates point features in multiple layers. We only take the second finest level of point features as a hidden state feature $\mathbf{H}^{A}$. We did not take the finest level of point cloud features, as they are redundant, as noted in previous work[41]. Utilizing the node indices maintained in ${}^{1\!}\mathcal{X}$, we can directly sample the node-wise points from $\mathbf{H}^{A}$. During the sampling, we only keep $K_{p}$ points for each node. If a node finds too few or too many points, we adjust the number to $K_{p}$ through random sampling or zero padding. The sampled node-wise points and features are $\mathcal{P}^{A}$ and $\mathbf{Z}^{A}$.
@@ VII Experiment
  • Self-supervised training. Our training approach, explained in Sec-VII-A4, involves data automatically generated from posed RGB-D sequences.
@@ VII-A2 Evaluation metrics
Node recall (NR) and node precision (NP) evaluate scene graph matching performance. If the intersection over the unit (IoU) between the nodes’ point cloud is higher than a threshold $\tau_{iou}=0.3$, we treat it as a true positive (TP) node pair. Otherwise, it is a false positive (FP) pair. Then, we can compute the NR and NP, similar to the recall and precision in image matching task [6]. In the registration task, we follow GeoTransformer[41] and use Inlier Ratio (IR) and Registration Recall (RR) to evaluate the accuracy. If the transformation error between the point cloud pair is within a threshold (i.e., RMSE$<$0.2m), we treat it as a successful registration. Registration recall is the portion of the successful registration.
@@ VII-C2 Registration Accuracy
To discuss the registration performance, we introduce the definition of pseudo inliers, the correspondences after outlier pruning. As introduced in Section-V, we prune the outliers in point correspondences via maximum clique. The remaining correspondences are called pseudo inliers. Similarly, GeoTransformer and SG-PGM prune outliers via a local-to-global registration [41], based on a weighted SVD[44]. We compute the pseudo inlier ratio (PIR), the ratio of true inliers within the pseudo inliers. As shown in Fig. 6, our IR and PIR are higher than those of GeoTransformer, demonstrating more accurate point matching. In the final registration result, both SG-Reg and GeoTransformer successfully align the scenes.
@@ VII-C3 Computational efficiency
SG-Reg requires significantly fewer floating-point operations per second (FLOPS) and consumes less GPU memory compared to baseline models. This enhanced computational efficiency arises from its ability to encode scenes at a higher compression ratio. As illustrated in Fig. 7(a), SG-Reg reduces a scene to dozens of semantic nodes, while GeoTransformer compresses it into hundreds of superpoints. Given that attentional operations [46] on an input feature set have a computational complexity of $O(n^{2})$, where $n$ is the number of input features, performing attentional operations on semantic nodes is far more computationally efficient than on superpoints. In the summary of computing resources presented in Fig. 7(b), SG-Reg consumes the least GPU memory and requires far fewer FLOPS than the two baseline models. Even for large scenes exceeding 210K points, our reserved GPU memory remains below 11GB, and the FLOPS are below $2,215$ MFLOPS. In contrast, GeoTransformer cannot inference the large scenes due to memory limitations; it exceeds the maximum GPU memory capacity of 24GB on the Nvidia RTX-3090. SG-PGM, our closest baseline, exhibits efficiency that falls between SG-Reg and GeoTransformer, as it inherits the superpoint encoding from GeoTransformer, necessitating more GPU resources than our approach.
@@ VII-E2 Evaluation metrics
With the predicted transformation $\mathbf{T}^{B}_{A}$ and ground-truth transformation $\bar{\mathbf{T}}^{B}_{A}$, we compute their relative translation error (RTE) and the relative rotation error (RRE). If RTE$<0.5m$ and RRE$<5^{\circ}$, we call it a successful registration.
@@ VII-E7 Qualitative results
Another interesting phenomenon is that our query frames with success registration are spatially close to each other, as shown in Fig 11(c). This is because our method requires a longer initialization. At the beginning of the sequence, our integrated scene graphs have very few semantic nodes and a low overlap ratio. As the scene graphs grow, they gradually have a larger intersection. Hence, nearly all of our query frames failed initially, while those in the latter half of the sequence were mostly successful. Conversely, as shown in Fig 11(b), the successful loop closure frames in HLoc are spatially distributed in variances.
To verify the stability of the three algorithms, we conducted a robust pose-averaging experiment to refine the predictions. According to our previous introduction, the robust pose average rejects outliers in relative pose estimation. Therefore, when the algorithm is robust enough to noise, the impact of robust pose average on accuracy will be minimal. When the algorithm has significant randomness, its accuracy may be significantly improved.
As shown in Fig. 11(e), pose averaging has a minimal impact on Hydra due to its low success rate. In contrast, HLoc achieves a higher success rate after pose averaging, demonstrating that HLoc is significantly affected by random noise. This also highlights the effectiveness of the pose averaging module in estimating a reliable pose from multiple predictions with variations. However, as shown in Fig. 11(f), pose averaging has little effect on our results. We attribute this to the inherent stability of our method: while it requires a relatively higher scene overlap and adopts a conservative approach to registration, it consistently achieves accurate and stable results without significant jitter.
@@
(a) Hydra
(b) HLoc
(c) Ours
Figure 11: Registration visualization from an office scenario. The first row shows the original results, while the second row shows the results after pose averaging using three nearest measurements. Each registration result is represented in a line. A registration is successful if its predicted camera pose has a relative translation error (RTE) of less than $0.2$m. For a better visualization, we align the camera poses with ground-truth transformation and incorporate a vertical translation
@@ VII-E8 Quantitative results
TABLE XI: Evaluate loop closure success rate and RTE. We report the average RTE from the successful frames. Pose Avg@$X$ refers to the robust pose average using $X$ frames.
Figure 12: The Loop closure success rate over the sequence length. Avg@$X$ refers to the robust pose average in $X$ consecutive loop frames.
@@
Meanwhile, we point out that SG-Reg is not adversarial with visual registration methods. In future works, they can be combined into a unified factor graph to optimize.
@@ VII-E9 Communication Bandwidth
To further evaluate the influence of communication rate and registration accuracy, we run an ablation study. It defines a minimum interval time for publishing dense messages. We change the interval time and evaluate the performance in one of the scanned scene.
@@
TABLE XIII: Ablation study of communication strategy. Group A turns off the dense message broadcast and fully relies on the coarse messages. BW refers to the averaged bandwidth over the query frames.
As shown in Table XIII, the success rate gradually increases as the interval time between publishing dense messages decreases. This demonstrates that dense messages significantly improve the likelihood of successful registration. However, the success rate approaches its optimal level at a certain communication rate, beyond which further improvements are marginal. In this experiment, we selected the setting of Group C for its optimal balance between registration success and communication efficiency.
@@
where $g(\cdot)$ is the similarity function in equation(12). In dense matching, it concatenates the point features in all the semantic nodes to construct $\{\tilde{\mathbf{Z}}^{A},\tilde{\mathbf{Z}}^{B}\}$ and run equation (20). They build the point features $\{\tilde{\mathbf{Z}}^{A},\tilde{\mathbf{Z}}^{B}\}$ differently, where we summarize their dimensions in Table XIV.
@@
CF ||×$|\mathcal{M}|\times K_{p}\times d_{z}$ ||×$|\mathcal{M}|\times K_{p}\times d_{z}$ $|\mathcal{M}|\cdot K_{p}\cdot K_{p}$
TABLE XIV: Analysis complexity in computing the similarity matrix. DM denotes dense matching, while CF denotes coarse-to-fine matching.
As shown in TABLEXIV, dense matching requires heavier dot product operations. To verify the analysis, we select a scene to perform dense matching. In inferencing the selected scene, the dense matching requires $36,182$ MFLOPS, while the coarse-to-fine matching requires $602$ MFLOPS. The recorded computational aligns with our theoretical analysis.
@@ VII-F2 Runtime in SLAM
However, although our key modules run faster than HLoc, we point out that our method can not run in real-time yet. One bottleneck is the point cloud pre-processing operation, adapted from GeoTransformer[41]. It uniformly downsamples the point cloud into multiple layers of point cloud. It searches the nearby points at each layer, generating a hierarchical point cloud aggregation structure. Due to the large size of our point cloud and it runs on CPU, it can take up to $900$ ms in one query frame. GeoTransformer and SG-PGM also run the pre-processing step and take a similar time. In the runtime summary, we treat the pre-processing operation as a step in the data loader. Because data loading time in HLoc is not involved either, we exclude the pre-processing time in TableXVI. Considering that real-time performance is not our motivation in this paper, we leave it to be better implemented in future work.
Hydra HLoc Ours
@@
TABLE XVI: (a) Registration time in the two-agent SLAM. We report the average time over the frame number. CPU running time is marked. (b) Semantic mapping time. FM-Fusion runs in real-time, while the foundation models run offline.

### v2

@@ I INTRODUCTION
Early works in semantic-aided registration explicitly maintain the semantic objects. To match re-observed semantic objects, they design hand-crafted descriptors and similarity metrics based on 2D image feature[9], projected bounding box [12, 8], quadratic representation[13], or graph topology[14, 10, 15]. Using the calculated similarity metric, they identify matched nodes by either setting a threshold or designing a set of verification rules. Although these works introduced pioneering SLAM frameworks that incorporate semantic landmarks, they often face long-tailed cases in real-world experiments due to the noise commonly present in the reconstructed semantic objects.
@@ II-C Scene Graph Construction
Fusion++[9] 2018 Explicit $\times$ $\times$ 2D Points Mask R-CNN -
@@
SG-Reg (Ours) 2024 Implicit Point Cloud 3D Points RAM-Ground-SAM NetVLAD+LightGlue
@@ II-D Semantic Data Association
Data association between semantic representations is the most important step to guarantee a successful registration. In semantic SLAM works, The step relies on explicit representations. Bowman et al. [12] represent each object using a bounding box. If two object nodes belong to the same semantic category and their intersection-over-union (IoU) exceeds a predefined threshold, they are considered a match. Building on Kimera, Hydra [15] constructs hand-crafted semantic descriptors by aggregating semantic histograms from nearby objects and computes node similarity using these descriptors. More recently, SlideSLAM [53] determines object similarity based on identical semantic categories and similar bounding box shapes. To incorporate topological information into semantic data association, X-View [14] introduces a random walk descriptor (RWD) that records semantic labels along walking routes. The descriptor represents the topology information explicitly. When calculating the similarity metric, X-View counts the number of rows with identical RWDs to determine a similarity score. To enhance the spatial descriptiveness of RWD, Lin et al. [10] integrate distance information into the arrangement of RWDs. Similarly, Liu et al. [55] utilize spatial priors to construct RWDs. To reduce false matches between nodes, some approaches [56] explicitly construct edge descriptors to verify geometric consistency among matched nodes. Along the same lines, Julia et al. [57] employ node triplets to validate the correctness of matched nodes.
@@
Inspired by the learning-based graph matching models [58], we believe training a neural network to learn semantic data association is a promising direction. The latest version of Hydra [50] proposes encoding object nodes using a Graph Neural Network (GNN), though it omits shape features. SGAligner [16] is the first work, to our knowledge, to focus on learning scene graph matching. It encodes multiple modalities from semantic nodes, including semantic labels, center positions, and relationship labels. Subsequently, SG-PGM enhances this approach by incorporating shape features into scene graph learning. It samples points from each semantic node and aggregates them using GATv2 [25]. SGAligner and SG-PGM adopted the geometric-related layers from GeoTransformer, including its superpoint matching, point matching, and local-to-global registration layer. Both of them are trained and evaluated on the 3RScan dataset [29], requiring ground-truth annotations to construct their scene graph.
@@ IV-B Triplet-boosted Graph Neural Network (GNN)
where the length of triplet edges $|\mathbf{e}_{i,j}|,|\mathbf{e}_{i,k}|\in\mathbb{R}$ and $\hat{\mathbf{e}}_{i,j},\hat{\mathbf{e}}_{i,k}$ are the normalized vectors. The term $(\hat{\mathbf{e}}_{i,j}\cdot\hat{\mathbf{e}}_{i,k})$ represents cosine value of the triplet angle. Sinusoidal functions $\psi^{L}$ and $\psi^{A}$ are applied to embed the edge length and the triplet angle. The sinusoidal functions have been used in GeoTransformer[41] to encode their geometric scalars.
@@ IV-C Shape Network
In the following steps, we explain the encoding for $\{{}^{l}\mathcal{X}\}_{l}$ and it goes the same way for $\{{}^{l}\mathcal{Y}\}_{l}$. As shown in Fig. 4, the point backbone generates point features in multiple layers. We only take the second finest level of point features as a hidden state feature $\mathbf{H}^{A}$. We did not take the finest level of point cloud features, because they are redundant, as noted in previous work[41]. Utilizing the node indices maintained in ${}^{1\!}\mathcal{X}$, we can directly sample the node-wise points from $\mathbf{H}^{A}$. During the sampling, we only keep $K_{p}$ points for each node. If a node finds too few or too many points, we adjust the number to $K_{p}$ through random sampling or zero padding. The sampled node-wise points and features are $\mathcal{P}^{A}$ and $\mathbf{Z}^{A}$.
@@ VII Experiment
  • Self-supervised training. Our training approach, explained in Sec-VII-A(4), involves data automatically generated from posed RGB-D sequences.
@@ VII-A2 Evaluation metrics
Node recall (NR) and node precision (NP) evaluate scene graph matching performance. If the intersection over the unit (IoU) between the nodes’ point cloud is higher than a threshold $\tau_{iou}=0.3$, we treat it as a true positive (TP) node pair. Otherwise, it is a false positive (FP) pair. Then, we can compute the NR and NP, similar to the recall and precision in image matching task [6]. In the registration task, we follow GeoTransformer[41] and use Inlier Ratio (IR) and Registration Recall (RR) to evaluate the accuracy. With the predicted transformation, if the root mean square error (RMSE) between the aligned point cloud is within a threshold (i.e., RMSE$<$0.2m), we treat the prediction as a successful registration. Registration recall is the portion of the successful registration.
@@ VII-C2 Registration Accuracy
To discuss the registration performance, we introduce the definition of pseudo inliers, the correspondences after outlier pruning and being treated as inliers by the estimator. As introduced in Section-V, we prune the outliers in point correspondences via maximum clique. Similarly, GeoTransformer and SG-PGM prune outliers via a local-to-global registration [41], based on a weighted SVD[44]. The baseline works and SG-Reg all have pseudo inliers. We compute the pseudo inlier ratio (PIR), the ratio of true inliers within the pseudo inliers. As shown in Fig. 6, our IR and PIR are higher than those of GeoTransformer, demonstrating more accurate point matching. In the final registration result, both SG-Reg and GeoTransformer successfully align the scenes.
@@ VII-C3 Computational efficiency
SG-Reg requires significantly fewer floating-point operations per second (FLOPS) and consumes less GPU memory compared to baseline models. This enhanced computational efficiency arises from its ability to encode scenes at a higher compression ratio. As illustrated in Fig. 7(a), SG-Reg reduces a scene to dozens of semantic nodes, while GeoTransformer compresses it into hundreds of superpoints. Given that attentional operations [46] on an input feature set have a computational complexity of $O(n^{2})$, where $n$ is the number of input features, performing attentional operations on semantic nodes is far more computationally efficient than on superpoints.
To verify our analysis, we design a study that runs a single GNN layer from GeoTransformer, SG-PGM and SG-Reg. We conduct the study in all of target scenes. As shown in in Fig. 7(b), SG-Reg consumes the least GPU memory and requires far fewer FLOPS than the two baseline models. Even for large scenes exceeding 210K points, our reserved GPU memory remains below 11GB, and the FLOPS are below $2,215$ MFLOPS. In contrast, GeoTransformer requires more than $1,000,000$ MFLOPS, significantly higher than ours. Besides, GeoTransformer cannot inference the large scenes due to memory limitations; it exceeds the maximum GPU memory capacity of 24GB on the Nvidia RTX-3090. SG-PGM, our closest baseline, exhibits efficiency that falls between SG-Reg and GeoTransformer, as it inherits the superpoint encoding from GeoTransformer, necessitating more GPU resources than our approach.
@@ VII-E2 Evaluation metrics
With the predicted transformation $\mathbf{T}^{B}_{A}$ and ground-truth transformation $\bar{\mathbf{T}}^{B}_{A}$, we compute their relative translation error (RTE) and the relative rotation error (RRE). If RTE$<0.2m$ and RRE$<5^{\circ}$, we call it a successful registration.
@@ VII-E7 Qualitative results
Original Refer to caption Refer to caption Refer to caption
Pose Avg@3 Refer to caption Refer to caption Refer to caption (a) Hydra (b) HLoc (c) SG-Reg (Ours)
Figure 11: Registration visualization in an office scenario. The first row shows the original results, while the second row shows the results after pose averaging using three nearest measurements. Each registration result is represented in a line. A registration is successful if its predicted camera pose has a $\text{RTE}<0.2m$ and $\text{RRE}<5^{\circ}$. For a better visualization, we align the camera poses with ground-truth transformation and incorporate a vertical translation
@@
Another interesting phenomenon is that our query frames with success registration are spatially close to each other, as shown in Fig 11. This is because our method requires a longer initialization. At the beginning of the sequence, our integrated scene graphs have very few semantic nodes and a low overlap ratio. As the scene graphs grow, they gradually have a larger intersection. Hence, nearly all of our query frames failed initially, while those in the latter half of the sequence were mostly successful. Conversely, as shown in Fig 11, the successful loop closure frames in HLoc are spatially distributed in variances.
To verify the stability of the three algorithms, we conducted a robust pose-averaging experiment to refine the predictions. According to our previous introduction, the pose averaging rejects outliers in the estimated transformation. Therefore, when the algorithm is robust enough to noise, pose averaging has a minimal impact. When the algorithm has drastic randomness, its accuracy can be improved after pose averaging.
As shown in Fig. 11, pose averaging has a minimal impact on Hydra due to its low success rate. In contrast, HLoc achieves a higher success rate after pose averaging, demonstrating that HLoc is significantly affected by random noise. This also highlights the effectiveness of the pose averaging module in estimating a reliable pose from multiple predictions with variations. However, as shown in Fig. 11, pose averaging has little effect on our results. We attribute this to the inherent stability of our method: while it requires a relatively higher scene overlap and adopts a conservative approach to registration, it consistently achieves accurate and stable results without significant jitter.
@@ VII-E8 Quantitative results
TABLE XI: Evaluate registration success rate (SR). We report the averaged RTE from the succeed registration. Pose Avg@$X$ refers to the robust pose average using $X$ frames.
Figure 12: The Loop closure success rate over the sequence length. Avg@$X$ refers to the robust pose averaging in $X$ consecutive loop frames.
@@
Meanwhile, we point out that SG-Reg is not adversarial with image registration methods. In future works, they can be combined into a unified factor graph to optimize.
@@ VII-E9 Communication Bandwidth
To further evaluate the influence of communication rate and registration accuracy, we run an ablation study. It defines a minimum interval time for publishing dense messages. We change the interval time and run the experiment in 4 settings. Results from each setting are summarized in Table. XIII.
@@
TABLE XIII: Ablation study of communication strategy. Setting A turns off the dense message broadcast and fully relies on the coarse messages. BW refers to the averaged bandwidth over the query frames.
As shown in Table XIII, from setting A to setting D, the success rate gradually increases as the interval time between publishing dense messages decreases. This demonstrates that dense messages significantly improve the likelihood of successful registration. However, the success rate approaches its optimal level at a certain communication rate, beyond which further improvements are marginal. In this experiment, we selected the setting of Group C for its optimal balance between registration success and communication efficiency.
@@
where $g(\cdot)$ is the similarity function in equation(12). In dense matching, it concatenates the point features in all the semantic nodes to construct $\{\tilde{\mathbf{Z}}^{A},\tilde{\mathbf{Z}}^{B}\}$ and run equation (20). They build the point features $\{\tilde{\mathbf{Z}}^{A},\tilde{\mathbf{Z}}^{B}\}$ differently, where we illustrate their computational complexity in Table XIV.
@@
CFM ||×$|\mathcal{M}|\times K_{p}\times d_{z}$ ||×$|\mathcal{M}|\times K_{p}\times d_{z}$ $|\mathcal{M}|\cdot K_{p}\cdot K_{p}$
TABLE XIV: Analysis complexity in computing the similarity matrix. DM denotes dense matching, while CFM denotes coarse-to-fine matching.
As shown in TABLE XIV, dense matching requires heavier dot product operations. To verify the analysis, we select a scene to perform dense matching. The result shows that the dense matching results in a high outlier ratio and requires higher computation FLOPS than the coarse-to-fine matching. Please check our supplementary material for details.
@@ VII-F2 Runtime in SLAM
However, although our key modules run faster than HLoc, we point out that our method can not run in real-time yet. One bottleneck is the point cloud pre-processing operation, adapted from GeoTransformer[41]. It uniformly downsamples the point cloud into multiple layers of point cloud. It searches the nearby points at each layer, generating a hierarchical point cloud aggregation structure. Due to the large size of our point cloud and it runs on CPU, it can take up to $900$ ms in one query frame. GeoTransformer and SG-PGM also run the same pre-processing step and take a similar time. Similar to them, we treat the pre-processing operation as a step in the data loader, which is not considered in the runtime. Besides, since data loading time in HLoc is not involved either, we exclude ours in TableXVI, which is a fair comparison. Considering that real-time performance is not our motivation in this paper, we leave it to be better implemented in future work.
Hydra HLoc SG-Reg
@@
TABLE XVI: (a) Registration time in the two-agent SLAM. We report the average time in each frame. Modules or method marked in run on CPU. (b) Semantic mapping time. FM-Fusion runs in real-time, while the foundation models run offline.